In a random sample of 100 audited estate tax​ returns, it was determined that the mean amount of additional tax owed was ​$3418 with a standard deviation of ​$2560. Construct and interpret a​ 90% confidence interval for the mean additional amount of tax owed for estate tax returns. What are the lower and upper bounds?
Answer #1

Solution :

Given that,

Point estimate = sample mean = = $3418

Population standard deviation =    = $2560

Sample size = n =100

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645( Using z table ( see the 0.05 value in standard normal (z) table corresponding z value is 1.645 )   


Margin of error = E = Z/2 * ( /n)

= 1.645* ( 2560 /  100 )

=421.12
At 90% confidence interval estimate of the population mean
is,

- E < < + E

3418 - 421.12   <   < 3418 + 421.12

2996.88 <   < 3839.12

lower bound 2996.88 and upper bound =3839.12

( 2996.88 ,3839.12 )At 90% confidence interval estimate of the population mean is between lower bound =2996.88 and upper bound =3839.12)